OpenStudy (anonymous):

Find the number of sides of a regular polygon if one interior angle is 120 degrees. A.) 6 B.) 5 C.) 4 D.) 3

OpenStudy (mathstudent55):

Do you know the formula that tells you the sum of the measures of the interior angles of a polygon?

OpenStudy (anonymous):

No

OpenStudy (mathstudent55):

Here is the formula: \(S = (n - 2)180\) where S = sum of the measures of the interior angles, and n = number of sides.

OpenStudy (mathstudent55):

Do you know what "regular" means when you say regular polygon?

OpenStudy (anonymous):

When all angles are equal?

OpenStudy (mathstudent55):

It means a polygon that has all sides of the same length and all interior angles of the same measure.

OpenStudy (mathstudent55):

Correct. All angles are congruent, and all sides are congruent.

OpenStudy (anonymous):

So would it be 3?

OpenStudy (mathstudent55):

Now we can solve this problem. Since you are dealing with a regular polygon, all angles have the same measure. Let's say the sum of the measures of the interior angles is S. Also, let's say your polygon has n sides. Then each angle would have the measure S/n Ok so far?

OpenStudy (mathstudent55):

No, it's not 3.

OpenStudy (mathstudent55):

Just follow what I am explaining . We'll find the answer soon.

OpenStudy (anonymous):

Okay!

OpenStudy (mathstudent55):

If the sum of the measures of the interior angles is S and the polygon has n sides, then each interior angle measures S/n.

OpenStudy (mathstudent55):

We have a formula for the sum of the measures of the interior angles. \(S = (n - 2)180\) If we divide it by n, the number of sides, we will have the measure of one angle. \(\dfrac{S}{n} = \dfrac{(n - 2)180}{n} \) is the measure of one angle. We are told one angle measures 120 deg, so we set our expression equal to 120 and ewe solve for n. \(\dfrac{(n - 2)180}{n} = 120\)

OpenStudy (anonymous):

I think I got it now! thank you so much!

OpenStudy (mathstudent55):

Now we have to solve that equation for n. We multiply both sides by n: \((n - 2)180 = 120n\) Distribute 180 on the left side: \(180n - 360 = 120n\) Subtract 120n from both sides, and add 360 to both sides: \(60n = 360\) Divide both sides by 60: \(n = 6\) Since n is the number of sides of the polygon, we can answer: The polygon has 6 sides.

OpenStudy (mathstudent55):

You're welcome.